### v1

# Characterisation of Open Quantum System Dynamics based on Information Back-flow

Abstract
For unital dynamics, we show that a generalized trace distance measure offers no advantage over the trace distance measure for witnessing non-Markovianity. We determine the class of non-unital channels where the standard trace distance measure is insufficient here and the generalized measure is necessary. Finally, we assess the status of the GTD measure as an indicator of information flow between an open system and its environment.
I Introduction
The isolation of a quantum system from its environment is an ideal condition, rendering open systems a significant object of study Breuer and Petruccione (2007). Such investigations are involved in achieving quantum control, focusing on the analysis of complete positive channels (CP) Choi (1975); Sudarshan et al. (1961). CP maps distinguish themselves from positive (P) maps Choi (1972) by remaining unaffected by including ancillary reference systems that do not participate during evolution. Any uncorrelated system and the environment with combined Hamiltonian evolution guarantees the CP evolution. Although certain quantum-correlated Shabani and Lidar (2009); Jagadish et al. (2023) states may induce CP dynamics, such occurrences are not universal across all unitary dynamics. The correlations generally lead to various dynamical maps, not just CP maps Jordan et al. (2004); Pechukas (1994); Štelmachovič and Bužek (2001); Colla et al. (2022); Dijkstra and Tanimura (2012).
From start to finish, the channel goes through different temporal dynamics depending on various possible environments and their interaction with the system. This time evolution determines different processes based on these interactions. Should the environment be sufficiently big or the interaction remains sufficiently weak, it is reasonable to assume that the environment does not vary too much alongside the system and that the correlations are not significant enough to be considered. Under such circumstances, the time evolution is laid out by the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) Gorini et al. (1976); Lindblad (1976) time-local equation, which follows the semi-group property characteristic of Markovian evolution. As quantum technology advances, allowing for experimentally greater control of the system’s dynamics, the above assumptions can be relaxed White et al. (2020); Bernardes et al. (2015); Gulati et al. (2024); Liu et al. (2013). In that situation, semi-group dynamics will be replaced by general divisible dynamics, provided the dynamics is invertible.
The divisibility of a dynamics is defined by:
$\mathcal{E}_{(t,0)}=\Lambda_{(t,\tau)}\circ\mathcal{E}_{(\tau,0)}$ (1)
where, $\Lambda$ is the intermediate map taking the evolution from time $\tau$ to $t$.
Semi-group dynamics is a special case of divisible dynamics where $\Lambda_{(t,\tau)}=\mathcal{E}_{t-\tau}$. In general, even if the dynamics $\mathcal{E}_{(t,0)}$ from the initial time to any final time $t$ is CP, the intermediate map may or may not be CP, depending on the system-environment interactions. Correspondingly, the divisible dynamics is categorized into CP-divisible and non-CP-divisible. Non-CP-divisible dynamics can be P-divisible and non-P-divisible, depending on whether the (non-CP) intermediate map is P or non-P. Non-CP-divisible dynamics is considered non-Markovian based on divisibility Rivas et al. (2010), while non-P-divisible dynamics is considered a stronger manifestation of non-Markovianity.
Breuer-Laine-Piilo Laine et al. (2010); Breuer et al. (2009) introduced an information-theoretic measure called BLP to understand the memory effect Breuer et al. (2016); Liu et al. (2013). It is based on the distinguishability of two quantum states, as the trace distance measure indicates. The CP and P processes weaken distinguishability, suggesting a loss of information to the environment. Since BLP does not respond to the translation part of the dynamics, a generalized trace distance measure (GBLP) was introduced. Here, the increment of generalized trace distance for finite time is identified with non-Markovianity. In an attempt to unify the non-Markovianity based on trace distance and divisibility, a theorem in Ref. Chruściński et al. (2018) shows that P-divisibility is connected to a trace distance condition, specifically the GBLP condition Chruściński et al. (2011); Wißmann et al. (2015) rather than the BLP condition. This would raise the hope that the GBLP measure should serve as the right indicator of the information flow Chruściński et al. (2018, 2011); Wißmann et al. (2015). However, here, we argue that it may not be so.
The paper is written in the following way: In II, the affine transformation of a qubit is defined, and the BLP and GBLP measures are introduced to characterize non-Markovianity. An example is shown for which BLP fails, but GBLP works. In Sec. III, the general class of dynamics where GBLP is necessary is determined. In Sec. IV, it is shown that GBLP is not the right indicator, in general, to capture the information flow. Finally, we conclude with discussions and conclusions in Sec. V.
II Non-Markovian measures for information backflow
An affine transformation can represent the dynamics of a qubit interacting with an arbitrary quantum environment. Let us write the state in Pauli basis, then $\rho=\dfrac{1}{2}(\mathbb{I}+\vec{r}\cdot\vec{\sigma})$ with $\lvert\vec{r}\rvert\leq 1$, here $\sigma^{\prime}s$ are Pauli matrices and $\vec{r}$ is a Bloch vector representing the state in the Bloch sphere. The Bloch vector transformation is given by
$\vec{r}^{\prime}=T\vec{r}+\vec{c}$ (2)
The transformation has two components: (a) a $3\times 3$ $T$ matrix representing the rotation, contraction and expansion, and (b) a $3\times 1$ $\vec{c}$ vector representing the translation. If $\vec{c}=\vec{0}$, then the dynamics is called unital dynamics, which leaves the completely mixed state, the centre of the Bloch sphere, invariant. The non-zero translation vectors represent non-unital dynamics.
II.1 BLP measure
Breuer-Laine-Piilo Laine et al. (2010); Breuer et al. (2009) defined a measure that connects non-Markovianity with the information back-flow. This measure is important because, through this, memory dependence on dynamics was highlighted, which was not possible through divisibility-based non-Markovianity. It is not necessary to know the dynamical map to witness the non-Markovian behaviour of the dynamics experimentally. For defining the measure, they considered the trace distance defined as
$D=\dfrac{1}{2}\lVert\rho_{1}-\rho_{2}\rVert_{1}$ (3)
here $\lVert A\rVert_{1}=\mathrm{Tr}\lvert A\rvert$, where modulus is defined as $\lvert A\rvert=\sqrt{A^{\dagger}A}$. This distance was shown to decide the distinguishability of two states prepared with equal probability:
$P_{\rm dis}=\dfrac{1+D}{2}$
here, $P_{\rm dis}$ is the probability to distinguish two states. Unitary dynamics keeps this distance unchanged for all initial pairs; hence, distinguishability is unchanged under unitary dynamics. This is consistent with the fact that the system does not interact with any other quantum system under unitary dynamics, so there is no exchange of information. In the event of interactions, based on distinguishability, there are two categories of dynamics: (a) Markovian dynamics: the distance does not increase for any initial pair for any interval of time. In this case, there is never an increase of distinguishability. (b) Non-Markovian dynamics: the distance increases at least for a few initial pairs for some time interval. In this case, the distinguishability may increase for an interval of time.
CP-divisible dynamics is CP for all intermediate-initial to intermediate-final times; this will keep contracting for all intermediate times since CP maps contract. P-divisible dynamics is CP from initial to final time, but for some intermediate time, it can be just P, not CP; it will also keep contracting because P-maps also contract. Because of this, CP-divisible and P-divisible dynamics are considered Markovian according to this measure Chruściński et al. (2018).
II.2 Generalized measure
As we can see from Eq. (3), this distance is unaffected by the translation vector $\vec{c}(t)$. So, the above measure will not capture any information flow associated with that.
Thus, the BLP definition does not respond to the translation part of the dynamics. If two maps have the same $T$ part with different $\vec{c}$, then BLP can not differentiate them. To overcome this, the GBLP measure, discussed below, was introduced. The Helstorm Matrix $\Delta_{p}(\rho_{1},\rho_{2})=p\rho_{1}-q\rho_{2}$ with $p+q=1$ was considered for the distance to overcome this issue because it is sensitive to translations Chruściński et al. (2011). The generalized distance is defined as
$D=\lVert\Delta_{p}(\rho_{1},\rho_{2})\rVert_{1},$ (4)
which is also a distance measure between two states but prepared with biased probabilities. This is also shown by connecting the distinguishability of two states with the generalized distance Chruściński et al. (2011); Wißmann et al. (2015). Unlike the BLP measure, P-divisibility is known to be connected with this generalised measure with the if and only if condition.
As a simple example where BLP fails but GBLP works, consider the non-P-divisible generalized amplitude damping noise given by the following affine map $\mathcal{E}(\vec{r})=T(t)\vec{r}+\vec{c}(t)$, where
$T(t)$ $=\begin{pmatrix}\sqrt{\eta}&0&0\\ 0&\sqrt{\eta}&0\\ 0&0&\eta\end{pmatrix}$
$\vec{c}(t)$ $=\begin{pmatrix}0\\ 0\\ (2s-1)(1-\eta)\end{pmatrix}$ (5)
where, $\eta=e^{-\gamma t}$ for $\gamma>0$ and $s=\cos^{2}{ft}$ for all $f$. Consider the initial states $\vec{r}_{1}\equiv(x_{1},y_{1},z_{1})^{\dagger}$ and $\vec{r}_{2}\equiv(x_{2},y_{2},z_{2})^{\dagger}$. We have the trace distance between their evolved versions to be
$\frac{1}{2}\textrm{Tr}[\mathcal{E}_{t}(\rho_{1})-\mathcal{E}_{t}(\rho_{2})]$ $=\frac{1}{2}|T(\textbf{r}_{1})-T(\textbf{r}_{2})|$
$=\frac{\sqrt{e^{-\gamma t}(\delta^{2}x+\delta^{2}y)+e^{-2\gamma t}\delta^{2}z}}{2},$ (6)
where $\delta x\equiv(x_{2}-x_{1})$, $\delta y\equiv(y_{2}-y_{1})$, and $\delta z\equiv(z_{2}-z_{1})$. The above equation implies a falling trace distance for all initial pairs, $\gamma$ and $f$, hence BLP Markovianity. With GBLP, the distance
$\textrm{Tr}[p\mathcal{E}_{t}(\rho_{1})-q\mathcal{E}_{t}(\rho_{2})]$ $\neq|pT(\vec{r}_{1})-qT(\vec{r}_{2})+(p-q)\vec{c}|$ (7)
As a simple illustration, for $\vec{r}_{1}\equiv(1,0,0)^{\dagger}$, $\vec{r}_{2}\equiv(0,1,0)^{\dagger}$, $p=0.25$, $f=4$, and $\gamma=0.1$, the required generalized trace distance is plotted in Fig (1). The distance is oscillating with time.
Figure 1: Generalised distance between states $\dfrac{1}{2}(\mathbb{I}+\sigma_{z})$ and $\dfrac{1}{2}(\mathbb{I}+\sigma_{x})$ for $p=0.25$ is plotted as a function of time for generalised amplitude damping noise (See Eq. II.2). Noise parameters $\gamma$ and $f$ are 0.1 and 4, respectively. Distance and time are in arbitrary units.
Thus, according to the theorem Chruściński et al. (2018), which states that non-P-visible dynamics will necessarily have non-decreasing generalized trace distance for some initial pair for a certain interval of time, this channel is non-P-divisible and hence non-CP-divisible. As such, it illustrates the usefulness of the GBLP measure where BLP fails.
III Unitality, divisibility and distance
The above represents one example where the two measures differ, but we want to determine the general classes where they differ. We are checking the distance evolution for both BLP and GBLP measures for CP, P and non-P maps. Once that is understood, the connection with the whole dynamics is straightforward because, for divisible dynamics, the whole dynamics is nothing but the composition of the intermediate maps. The GBLP measure is calculated for two initial states as $\rho_{i}\equiv(x_{i},y_{i},z_{i})$, where $i=1,2$ and probability $p$. The eigenvalues of $\Delta^{\dagger}\Delta$ are
$\dfrac{1}{4}\bigg{(}\sqrt{(px_{1}-qx_{2})^{2}+(py_{1}-qy_{2})^{2}+(pz_{1}-qz_{2})^{2}}\pm\lvert p-q\rvert\bigg{)}^{2}$
To calculate the distance, we add the square root of the eigenvalues because the trace does not depend on the basis. The generalised distance is
$D=\dfrac{1}{2}\lvert\lvert\vec{w}\rvert+\lvert p-q\rvert\rvert+\lvert|\vec{w}\rvert-\lvert p-q\rvert\rvert$
Here, $\vec{w}=p\vec{r_{1}}-q\vec{r_{2}}$. After expanding the above equation,
$D(t)=\begin{array}[]{ll}\lvert\vec{w}(t)\rvert,&\lvert\vec{w}(t)\rvert\geq\lvert p-q\rvert\\ p-q,&\lvert\vec{w}(t)\rvert<\lvert p-q\rvert\\ \end{array}$ (8)
We notice that if $p=q$, the GBLP distance reverts to the BLP distance.
Theorem 1.
The GBLP and BLP conditions are equivalent for unital dynamics.
Proof.
The translation vector vanishes for unital dynamics, and the dynamics is characterized by rotation, compression or expansion. Eq. (8) reduces to:
$D(t)=\begin{array}[]{ll}\lvert T(p\vec{r}_{1}-q\vec{r}_{2})\rvert,&\lvert\vec{w}(t)\rvert\geq\lvert p-q\rvert\\ p-q,&\lvert\vec{w}(t)\rvert<\lvert p-q\rvert\\ \end{array}$ (9)
We note the following:
  1. 1.
    If the channel is GBLP NM, then $\lvert T(p\vec{r}_{1}-q\vec{r}_{2})\rvert>\lvert p\vec{r}_{1}-q\vec{r}_{2}\rvert$ for some $p$, $\vec{r}_{1}$, $\vec{r}_{2}$, and $t$. Since $p,q\leq 1$, $p\vec{r}_{1}$ and $q\vec{r}_{2}$ will also be valid vectors belonging to the Bloch sphere. Hence $\lvert T(\vec{r^{\prime}}_{1}-\vec{r^{\prime}}_{2})\rvert>\lvert\vec{r^{\prime}}_{1}-\vec{r^{\prime}}_{2}\rvert$ where $\vec{r^{\prime}}_{1}=p\vec{r}_{1}$ and $\vec{r^{\prime}}_{2}=q\vec{r}_{2}$. This is the condition for BLP-NM.
  2. 2.
    If BLP is NM, $\lvert T(\vec{r}_{1}-\vec{r}_{2})\rvert>\lvert\vec{r}_{1}-\vec{r}_{2}\rvert$ for some $\vec{r}_{1}$ and $\vec{r}_{2}$. This condition gives one pair that belongs to GBLP with $p=\dfrac{1}{2}=q$. That makes GBLP also NM.
As a corollary of Theorem 1, there is no unital non-P-divisible BLP Markovian dynamics. We now consider the question of what are the channels for which the BLP and GBLP criteria are inequivalent.
Theorem 2.
BLP and GBLP differ only for non-unital non-P-divisible BLP dynamics.
Proof.
Our earlier example showed that the BLP and GBLP conditions are inequivalent. From Theorem 1, it follows that, nevertheless, they are equivalent for unital noise. Thus, their inequivalence can hold only for non-unital noise. Here, by Ref. Chruściński et al. (2018), we know that P-divisibility is equivalent to the GBLP Markovianity. Thus, the inequivalence of the BLP and GBLP conditions can arise only from non-unital non-P-divisible noise. If this noise is BLP non-Markovian, then the two criteria will agree. ∎
IV Information flow and distinguishability
The idea of information flow was introduced as an interpretation of distinguishability in the BLP definition. If the distinguishability is invariant between all pairs, then this shows no information exchange, indicating that the dynamics is unitary. On the other hand, if the distance monotonically decreases, the distinguishability decreases, suggesting that the information should be transferred away from the system.
This intuition was justified by defining the internal information ($I_{int}$) between the system states by BLP and GBLP distances, which can be accessed by the measurement of the system alone Breuer et al. (2016). The external information ($I_{ext}$) which cannot be accessed alone from the system lies in correlation or/and within environment states and is defined by subtracting from total information ($I_{total}$), which is from the combined state of the system-environment. So
$I_{int}^{t}(\rho_{S}^{1},\rho_{S}^{2})+I_{ext}^{t}=I_{total}^{t}(\rho^{1},\rho^{2}),$ (10)
where $I_{ext}=I_{total}-I_{int}$ and $\rho_{S}$, $\rho$ are system and combined system-environment state respectively. Since total information is constant owing to the unitarity of the dynamics, $I_{total}^{t}=I_{total}^{0}$. All arguments of information forward-flow or back-flow are justified by Eq. (10). We now consider how well the shift behaviour in a non-unital noise connects with the distinguishability condition.
IV.1 False positive test of distinguishability
To give an example of a case where the GBLP condition can be misleading, consider the dynamical map with $T=\left(\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\\ \end{array}\right)$, which takes all the states to the centre of the Bloch sphere, followed by shift $\vec{c}=\left(\begin{array}[]{c}0\\ 0\\ c\end{array}\right)$ which takes that centre to $c$. If we apply the GBLP measure, the distance will be $\lvert(p-q)\rvert$ or $\lvert(p-q)c\rvert$ depending on the value of $c$, which is non-zero. Though the states are identical, this distance gives the appearance of distinguishing them because it has a label of biased probabilities, which reflects the preparation information rather than information back-flow.
The divisible dynamics mentioned in theorem 2 where GBLP is necessary will always give false positives. That can specifically be seen in the example shown in II.2.
IV.2 False negative test of distinguishability
The GBLP definition does not follow distinguishability and information flow for all pairs. There is a possibility states can evolve and come closer or go far, but the information flow is zero, which misguides the interpretation of distinguishability and information flow.
Examples: Let us consider the unital dynamics with
$T=\left(\begin{array}[]{ccc}\exp^{-\gamma t}&0&0\\ 0&\exp^{-\gamma t}&0\\ 0&0&\exp^{-\gamma t}\end{array}\right)$ (11)
According to Eq. (10), the BLP definition suggests the information outflow for all pairs for all time. In light of Eq. (9), the generalized distance is constant for Bloch radius $r\leq 0.5$ in the considered example for $\gamma=0.1$ and $p=0.25$. Thus, GBLP will show no outflow or distinguishability change for certain pairs for corresponding probabilities, as shown in Fig (2). This goes against our expectation that the absence of information flow should only be true for unitary dynamics or the pairs where their sub-spaces do not evolve. This suggests that the generalized definition of distance does not capture the information flow for all pairs as required by Eq. 10.
Figure 2: Generalised distance is plotted as a function of time for different Bloch radii $r$ with two diametrically opposite points for an isometrically decaying dynamical map (Eq. 11). The decay parameter $\gamma$ and the probability of preparing states $p$ are 0.1 and 0.25, respectively. Distance and time are in arbitrary units.
Figure 3: Generalised distance is plotted as a function of time for different pairs of diametrically opposite points on the surface of Bloch sphere for a spin interacting with another spin (Eq. 12). The dynamical parameter $\omega$ and the probability of preparing states $p$ are 1.25 and 0.25, respectively. Distance and time are in arbitrary units.
This issue is not restricted to BLP Markovian dynamics. For BLP non-Markovian dynamics, the domain exists, showing no outflow or inflow, even though states manifestly evolved during that period. To show this, two interacting spins are considered; one spin is considered a system and the other an environment. The dynamical map is given by:
$T=\left(\begin{array}[]{ccc}\cos{\omega t}&0&0\\ 0&\cos{\omega t}&0\\ 0&0&1\\ \end{array}\right)$ (12)
This non-invertible process is trivially non-Markovian and has an evident information back-flow. Yet, for certain pairs of states, it can evolve their distance from a regime of $\lvert\vec{w}(t)\rvert\geq\lvert p-q\rvert$ to one of $\lvert\vec{w}(t)\rvert<\lvert p-q\rvert$, which in light of Eq. (8), can cause the GBLP measure to flatten out. The non-changing phase is shown in Fig (3) for $\omega=1.25$ and $p=0.25$, which shows no information flow from or towards the system. We can notice in this example that pairs taken on the z-axis will not evolve, which is consistent for both BLP and GBLP measures.
These examples suggest that GBLP is a measure of contracting maps related to P-divisibility. This does not capture the idea of information flow for all cases.
V Discussion and Conclusion
The generalised BLP measure was needed to characterize non-Markovianity because BLP did not capture aspects of the non-unital noise. In specific, it was shown that GBLP has a direct connection with P-divisibility. We have demonstrated that GBLP and BLP are equivalent witnesses of non-Markovianity for unital dynamics. For non-unital cases, where the Bloch sphere is contracting, BLP does not help decide the non-Markovianity. In those cases, GBLP shows that distance can increase with time solely because of translation, which is associated with information back-flow. Here, we have indicated certain situations that undermine this association. The GBLP condition can yield a false positive result as an artifact of initial biased preparation. Conversely, we also indicate instances where it yields a false negative result when some initial pairs do not change the generalized distance despite the evident effect of the environment on their subsystem dynamics. Thus, the GBLP condition is only correct for capturing the divisibility properties, but not information flow for all cases.
Acknowledgments
VP acknowledges the financial assistance of the Indian Science and Engineering Research Board (SERB) through grant CRG/2022/008345.